OpenStudy (crashonce):

A sold an article to B at a profit of 30%. B sold it to C for $2912 at a loss of 20%. THe original cost of the article is?

OpenStudy (crashonce):

@mathmath333

OpenStudy (mathmath333):



OpenStudy (crashonce):

then?

OpenStudy (crashonce):

oh

OpenStudy (crashonce):

is it $2800 thanks

OpenStudy (mathmath333):

this will make it clear \(\large\tt \begin{align} \color{black}{x\implies x+30\%x\\=x+\dfrac{30x}{100}\\~\\ let y=x+\dfrac{30x}{100}\\~\\ =\dfrac{130}{100}x\\~\\ y\implies y-\%20y \\ =y-\dfrac{20}{100}y\\ =\dfrac{80}{100}y\\ =\dfrac{80}{100}\dfrac{130}{100}x\\~\\~\\~\\ \dfrac{80}{100}\dfrac{130}{100}x=2192 }\end{align}\)

OpenStudy (mathmath333):

* correction last line\(\large\tt \begin{align} \color{black}{ \dfrac{80}{100}\dfrac{130}{100}x=2912 }\end{align}\)

OpenStudy (triciaal):

one approach let the original price of the article = P then 2 equations B paid 30% more = 1.3P and C paid 2912 at 20%loss so 80% of (1.3P) 0.80*(1.3P) = 2912 to solve for P 2912/ (0.80*1.3) = 2800